ROTATIONAL SPECTRA OF N_{2}OH^{+} AND CH_{2}CHCNH^{+} MOLECULAR IONS

\underline{OSCAR MARTINEZ, JR.}, VALERIO LATTANZI, and MICHAEL C. McCARTHY, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and School of Engineering and Applied Science, Harvard University, Cambridge, MA 02138; SVEN THORWITH, Max-Planck-Institut für Radioastronomie, Bonn, Germany, and I. Physikalisches Institut, Universität zu Köln, Germany.

Protonated molecular ions of nitrous oxide (N_{2}OH^{+}) and acrylonitrile (CH_{2}CHCNH^{+}) have been detected at high spectral resolution in the molecular beam of a Fourier transform microwave spectrometer on the basis of high-level ab initio calculations. The ions were synthesized in the throat of a pulsed supersonic nozzle by discharging in a flow of the corresponding precursor gas (either N_{2}O or CH_{2}CHCN) heavily diluted in H_{2}. Two isomers of N_{2}OH^{+} were identified, corresponding to protonation at either the N or O end of NNO. This work contributes precise nitrogen hyperfine coupling constants to existing measurements of ground state NNOH^{+}, and represents the first detection of the higher energy HNNO^{+} isomer, which is calculated to lie 4.4 kcal/mol above ground^{a}. In addition, protonated acrylonitrile has been detected for the first time at high spectral resolution, yielding spectroscopic constants that are in excellent agreement with high-level quantum-chemical calculations^{b}. Owing to sizable calculated dipole moments of protonated nitrous oxide and acrylonitrile and the relatively high proton affinities of their neutral counterparts, both cations are plausible candidates for astronomical detection with radio telescopes.

aJ. M. L. Martin & T. J. Lee, J. Chem. Phys. 98, 7951 (1993)

bAt CCSD(T)/cc-pwCVQZ level of theory with zero-point vibrational effects at CCSD(T)/cc-pVQZ.